# Suppression of the type Ia supernova host galaxy step in the outer regions of galaxies

### v1

@@ Abstract
Using 1533 type Ia supernovae (SNe Ia) from the five-year sample of the Dark Energy Survey (DES), we investigate the effects of projected galactocentric separation between the SNe and their host galaxies on their light curves and standardization. We show, for the first time, that the difference in SN Ia post-standardization brightnesses between high and low-mass hosts reduces from $0.078\pm 0.011$ mag in the full sample to $0.036\pm 0.018$ mag for SNe Ia located in the outer regions of their host galaxies, while increasing to $0.100\pm 0.014$ mag for SNe in the inner regions. In these inner regions, the step can be reduced (but not removed) using a model where the $R_{V}$ of dust along the line-of-sight to the SN changes as a function of galaxy properties. To explain the remaining difference, we use the distributions of the SN Ia stretch parameter to test whether the inferred age of SN progenitors are more varied in the inner regions of galaxies. We find that the proportion of high-stretch SNe Ia in red (older) environments is more prominent in outer regions and that the outer regions stretch distributions are overall more homogeneous compared to inner regions, but conclude that this effect cannot explain the reduction in significance of any Hubble residual step in outer regions. We conclude that the standardized distances of SNe Ia located in the outer regions of galaxies are less affected by their global host galaxy properties than those in the inner regions.
@@
pubyear: 2024pagerange: Suppression of the type Ia supernova host galaxy step in the outer regions of galaxiesSuppression of the type Ia supernova host galaxy step in the outer regions of galaxies
@@ 1 Introduction
Type Ia supernovae (SNe Ia) are a mature cosmological probe. The industrialisation of their discovery, measurement and analysis provides a ‘standardizable candle’ giving unparalleled insight into the Universe’s expansion history, particularly in the redshift $z<0.6$ Universe where dark energy dominates. However, despite the established consensus of an exploding carbon-oxygen white dwarf star with a light curve powered by the radioactive decay of 56Ni, there remain open questions over the details of the configuration of the progenitor systems, their explosion physics, and the effect of dust on their light curves and luminosities.
It is well known that SN Ia photometric properties have dependencies on the environment in which the SNe explode. For example, intrinsically brighter and slower evolving SNe Ia tend to explode in younger stellar populations (Hamuy et al., 2000). For cosmological applications, such large scale environmental effects can be removed after the application of standardization relations between SN Ia luminosity and SN light-curve width (Phillips, 1993), and SN Ia luminosity and SN optical colour (Riess et al., 1996; Tripp, 1998).
More concerning for cosmological measurements is the observation that, even after such standardization, the brightnesses of SNe Ia (and thus the distances inferred to them) have a residual dependence on the properties of the galaxy in which they exploded. SNe Ia in massive, passive, older galaxies are brighter post standardization, and those in low-mass, younger, star-forming galaxies are fainter. Simplistically, this manifests as a step in SN Ia post-standardization luminosity at a host galaxy stellar mass close to ${}^{10}\,\mathrm{M}_{\sun}$ (Kelly et al., 2010; Sullivan et al., 2010; Lampeitl et al., 2010). This so-called ‘mass step’ is observed or routinely modelled in all large SN Ia surveys or compilations of datasets (e.g., Betoule et al., 2014; Brout et al., 2022; DES Collaboration et al., 2024; Ginolin et al., 2024a) and has a typical size of about 0.06–0.15 mag (or three to seven per cent in distance) depending on the details of the sample. Similar trends are seen when replacing host stellar mass with other global host galaxy properties such as star-formation rate (SFR; Rigault et al., 2013), rest-frame colours (Roman et al., 2018; Jones et al., 2018b) or gas-phase/stellar metallicity (Childress et al., 2013; Moreno-Raya et al., 2016; Millán-Irigoyen et al., 2022), and when considering these properties measured locally in a small aperture at the SN position (Rigault et al., 2013; Roman et al., 2018; Kelsey et al., 2021).
@@
We use the Dark Energy Survey (DES) SN programme five-year dataset (Sánchez et al., 2024) to examine the effect of galactocentric distance on SN Ia standardization. The DES-SN5YR dataset is a well-calibrated and understood high-redshift SN Ia cosmological dataset, with deep imaging data available for the SN Ia host galaxies (Wiseman et al., 2020). Throughout, where relevant we assume a reference cosmology of a flat $\Lambda$CDM universe with $\Omega_{\mathrm{matter}}=0.3$, and with a Hubble constant of $H_{0}=70$ km s-1 Mpc-1.
@@ 2 Data
The host galaxies themselves were identified from a catalogue generated from deep, stacked $griz$ images from DES free of SN light (Wiseman et al., 2020), using the Source Extractor software (Bertin & Arnouts, 1996) to measure host positions, structural parameters and photometry. SNe were matched to candidate host galaxies using the directional light radius (DLR) method (Sullivan et al., 2006; Gupta et al., 2016; Sako et al., 2018). The host galaxy photometry has been fit with a series of galaxy spectral energy distribution (SED) templates to estimate host stellar masses ($M_{*}$), SFRs and $U-R$ rest-frame colours following Sullivan et al. (2006) and Kelsey et al. (2021) as implemented in DES-SN5YR (Smith et al., 2020). Galaxy rest-frame colours are reported in Bessell (1990) passbands and in the Vega magnitude system (Johnson & Morgan, 1953). These rest-frame colours are calculated using the method outlined in Kelsey et al. (2021), where the best-fitting host galaxy SED for each SN is adjusted with a wavelength-dependent multiplicative function such that the SED exactly reproduces the observed DES $griz$ photometry (often referred to as ‘mangling’ Hsiao et al., 2007; Conley et al., 2008).
@@ 2.1 SN Ia distances and bias corrections
where $m_{x}$, $x_{1}$ and $c$ are the SALT3 SN light-curve parameters, $\alpha$ and $\beta$ are global nuisance parameters parametrizing the stretch–luminosity and colour–luminosity relations, and M is the absolute magnitude of a SN Ia with $x_{1}=0$ and $c=0$. $\mu_{\mathrm{bias}}$ is the bias term that corrects for various observational and astrophysical selection effects.
@@
where $\mu_{\mathrm{obs}}$ is defined in equation 1 and $\mu_{\mathrm{model}}$ is the distance modulus in a reference cosmology. The Hubble residuals indicate whether a given SN Ia is brighter or fainter than expected in the reference cosmology; negative residuals indicate a brighter event.
In our analysis we differ from the main DES-SN5YR analysis in three ways. Firstly, we do not include a ‘$\gamma G_{\mathrm{host}}$’ term in equation 1. When used, such a term is designed to encapsulate any residual dependencies between SN Ia luminosities and their host galaxy properties (such as a stellar mass step) that are not corrected for in the light-curve fitting or bias correction steps. As a goal of our paper is to investigate such host galaxy effects, we do not apply this correction.
@@
In estimating our distances using BBC 1D, we recover nuisance parameter values of $\alpha=0.159\pm 0.009$ and $\beta=2.73\pm 0.05$. These differ from those reported in table 5 of Vincenzi et al. (2024), as those values were estimated using the BBC 4D bias corrections, but are similar to other estimates of $\alpha$ and $\beta$ found for the DES-SN5YR sample using 1D corrections (e.g., Kelsey et al., 2023).
@@ 2.2 SN–host separations
SN celestial coordinates ($x_{\mathrm{SN}}$, $y_{\mathrm{SN}}$) were measured by DES using point-spread function (PSF) fitting photometry (Kessler et al., 2015), with the coordinates computed using the weighted-average from each measurement during an observing ‘season’. Host galaxies were detected and measured in the DES image stacks by (Wiseman et al., 2020) using Source Extractor. We use the celestial coordinates (image centroids) of the Source Extractor detections ($x_{\mathrm{host}}$, $y_{\mathrm{host}}$) measured from images with the same world coordinate system as the images on which $x_{\mathrm{SN}}$, $y_{\mathrm{SN}}$ are determined. These then define the angular separation (or projected galactocentric distance), $\Delta\theta$, between the SN and the centre of the host galaxy, i.e.,
@@
To measure DLRs, we use the Source Extractor basic shape parameters A (semi-major axis) and B (semi-minor axis) that define the elliptical shape of the detected galaxy. The scales of A and B are set by the second-order moments (variance) of the galaxy’s profile along the A and B axes (e.g., Stobie, 1980).
@@ 3.2.1 Hubble residuals as a function of $d_{\mathrm{DLR}}$
The trend itself is not unexpected given the known relationships between Hubble residual and stellar mass (e.g., left panel of Fig. 4): regions within $d_{\mathrm{DLR}}<0.5$ are likely to be similar to the old, passive stellar environments of massive galaxies, and the outer regions similar to those in younger lower mass galaxies. A logical next step is to test how the Hubble residual–$d_{\mathrm{DLR}}$ trend relates to the well-known stellar mass and host galaxy colour steps.
@@ 3.2.2 Host stellar mass step and $d_{\mathrm{DLR}}$
We show the host galaxy stellar mass step separately for SNe Ia in bins of $d_{\mathrm{DLR}}$ and find that the size of the stellar mass step does depend on the galactocentric distance (Fig. 5). Step sizes and their significance are reported in Table 2. The greatest disparity occurs at $d_{\mathrm{DLR}}=1$: within that radius, SNe have a mass step of $0.100\pm 0.014$ mag ($6.9\,\sigma$), whereas for $d_{\mathrm{DLR}}>1$ the step is consistent with zero ($0.036\pm 0.018$ mag; $2.0\,\sigma$). Recall that we have made no attempt in our bias correction methodology to model the host mass step: the step is simply not present in our sample with $d_{\mathrm{DLR}}>1$.
@@
Sample
Number
(LM hosts, HM hosts)
Mean $\Delta\mu$ in
low-mass hosts
Mean $\Delta\mu$ in
high-mass hosts
Size of the
mass step
Significance
of step
r.m.s in
low-mass hosts
r.m.s in
high-mass hosts
@@
$\mathrm{d_{DLR}}\leq 1$ 1028 (310, 718) $0.065\pm 0.012$ $-0.035\pm 0.009$ $0.100\pm 0.014$ $6.9\sigma$ 0.205 0.232
$\mathrm{d_{DLR}}>1$ 505 (162, 343) $0.023\pm 0.014$ $-0.014\pm 0.011$ $0.036\pm 0.018$ $2.0\sigma$ 0.180 0.212
@@
To determine if the observed difference in mass step size between inner and outer regions is best fit by a single step at $d_{\mathrm{DLR}}=1$, we use smaller $d_{\mathrm{DLR}}$ bins, each with a similar number of SNe Ia. We also fit a straight line to the data. The results are shown in Fig. 7. The figure shows a clear difference in step size between $d_{\mathrm{DLR}}\leq 1$ and $d_{\mathrm{DLR}}>1$, i.e., the result in Fig. 5 is not dependent on the choice of $d_{\mathrm{DLR}}=1$ to divide the sample.
@@
Figure 8: As Fig. 5 but for host galaxy rest-frame $U-R$ colour. The step sizes are $0.106\pm 0.013$ mag (left) and $0.039\pm 0.018$ mag (right).

### v2

# Reduction of the type Ia supernova host galaxy step in the outer regions of galaxies

@@ Abstract
Using 1533 type Ia supernovae (SNe Ia) from the five-year sample of the Dark Energy Survey (DES), we investigate the relationship between the projected galactocentric separation of the SNe and their host galaxies and their light curves and standardization. We show, for the first time, that the difference in SN Ia post-standardization brightnesses between high and low-mass hosts reduces from $0.078\pm 0.011$ mag in the full sample to $0.036\pm 0.018$ mag for SNe Ia located in the outer regions of their host galaxies, while increasing to $0.100\pm 0.014$ mag for SNe in the inner regions. The difference in the size of the mass step between inner and outer regions is $0.064\pm 0.023$ mag. In these inner regions, the step can be reduced (but not removed) using a model where the $R_{V}$ of dust along the line-of-sight to the SN changes as a function of galaxy properties. We investigate the remaining difference using the distributions of the SN Ia stretch parameter to test the inferred age of SN progenitors. Comparing red (older) environments only, outer regions have a higher proportion of high-stretch SNe and a more homogeneous stretch distribution. However, this effect cannot explain the reduction in significance of any Hubble residual step in outer regions. We conclude that the standardized distances of SNe Ia located in the outer regions of galaxies are less affected by their global host galaxy properties than those in the inner regions.
@@
pubyear: 2025pagerange: Reduction of the type Ia supernova host galaxy step in the outer regions of galaxiesA
@@ 1 Introduction
Type Ia supernovae (SNe Ia) are a mature cosmological probe. The industrialisation of their discovery, measurement and analysis provides a ‘standardizable candle’ giving unparalleled insight into the Universe’s expansion history, particularly in the redshift $z<0.6$ Universe where dark energy dominates. However, despite the established consensus of an exploding carbon-oxygen white dwarf star with a light curve powered by the radioactive decay of 56Ni, there remain open questions over the details of the configuration of the progenitor systems, their explosion physics, and the effect of dust on their light curves and luminosities (for reviews, see Maguire, 2016; Liu et al., 2023).
For cosmological applications, SN Ia brightnesses are standardized by the application of relations between SN Ia luminosity and SN light-curve width (Phillips, 1993), and SN Ia luminosity and SN optical colour (Riess et al., 1996; Tripp, 1998). It is well known that some SN Ia photometric properties used in this standardization have dependencies on the environment in which the SNe explode. For example, intrinsically brighter and slower evolving SNe Ia tend to explode in younger stellar populations (Hamuy et al., 2000).
More concerning for cosmological measurements is the observation that, even after such standardization, the brightnesses of SNe Ia (and thus the distances inferred to them) have a residual dependence on the properties of the galaxy in which they exploded. SNe Ia in massive, passive, older galaxies are brighter post standardization, and those in low-mass, younger, star-forming galaxies are fainter. Simplistically, this manifests as a step in SN Ia post-standardization luminosity at a host galaxy stellar mass close to ${}^{10}\,\mathrm{M}_{\sun}$ (Kelly et al., 2010; Sullivan et al., 2010; Lampeitl et al., 2010). This so-called ‘mass step’ is observed or routinely modelled in all large SN Ia surveys or compilations of datasets (e.g., Betoule et al., 2014; Brout et al., 2022; Rubin et al., 2023; DES Collaboration et al., 2024; Ginolin et al., 2024a) and has a typical size of about 0.06–0.15 mag (or three to seven per cent in distance) depending on the details of the sample. Similar trends are seen when replacing host stellar mass with other global host galaxy properties such as star-formation rate (SFR; Rigault et al., 2013), rest-frame colours (Roman et al., 2018; Jones et al., 2018b) or gas-phase/stellar metallicity (Childress et al., 2013; Moreno-Raya et al., 2016; Millán-Irigoyen et al., 2022), and when considering these properties measured locally in a small aperture at the SN position (Rigault et al., 2013; Roman et al., 2018; Kelsey et al., 2021).
@@
A complementary approach to investigating the mass step is to measure the effect that alternative standardisation procedures have on its size and significance. Examples of this include using near-infrared peak magnitudes (Uddin et al., 2020; Ponder et al., 2021; Johansson et al., 2021), Bayesian hierarchical models (e.g. Thorp & Mandel, 2022), or the spectral modelling of ‘twin’ SNe Ia (Boone et al., 2021). Some of these approaches have led to a reduced mass step. Johansson et al. (2021) found a mass step consistent with zero after allowing $R_{V}$ to vary between each SN in their light curve fit, although Thorp & Mandel (2022) found that the step persisted regardless of their treatment of $R_{V}$. Boone et al. (2021) reduced the step from $0.092\pm 0.024$ to $0.040\pm 0.020$ when standardizing using a twins embedding approach. While promising in understanding the underlying cause of the step these methods all necessitate either expensive data and/or low-$z$ SNe Ia and cannot be used in the standardization of the current state-of-the-art large high-$z$ analyses.
@@
We use the Dark Energy Survey (DES) SN programme five-year dataset (Sánchez et al., 2024) to examine the relationship between galactocentric distance and SN Ia standardization. The DES-SN5YR dataset is a well-calibrated and understood high-redshift SN Ia cosmological dataset, with deep imaging data available for the SN Ia host galaxies (Wiseman et al., 2020). Throughout, where relevant we assume a reference cosmology of a flat $\Lambda$CDM universe with $\Omega_{\mathrm{matter}}=0.3$, and with a Hubble constant of $H_{0}=70$ km s-1 Mpc-1.
@@ 2 Data
Contamination in our SN Ia sample from non-SNIa events has been extensively considered and modelled in previous articles. The DES approach used an improved set of spectrophotomteric core collapse SN templates (Vincenzi et al., 2019), with expected level of contamination in cosmological samples studied in (Vincenzi et al., 2021) and (Vincenzi et al., 2023). The level of contamination based on these simulations is expected to be 0.8–3.2 per cent. On the real DES data, different classification algorithms also lead to variations in the SN Ia numbers at a similar level (fig. 11 in Vincenzi et al., 2024).
The host galaxies were identified from a catalogue generated from deep, stacked $griz$ images from DES free of SN light (Wiseman et al., 2020), using the Source Extractor software (Bertin & Arnouts, 1996) to measure host positions, structural parameters, and photometry using Kron (1980)-like apertures. SNe were matched to candidate host galaxies using the directional light radius (DLR) method (Sullivan et al., 2006; Gupta et al., 2016; Sako et al., 2018). The host galaxy photometry has been fit with a series of galaxy spectral energy distribution (SED) templates to estimate host stellar masses ($M_{*}$), SFRs and $U-R$ rest-frame colours following Sullivan et al. (2006) and Kelsey et al. (2021) as implemented in DES-SN5YR (Smith et al., 2020). Galaxy rest-frame colours are reported in Bessell (1990) passbands and in the Vega magnitude system (Johnson & Morgan, 1953). These rest-frame colours are calculated using the method outlined in Kelsey et al. (2021), where the best-fitting host galaxy SED for each SN is adjusted with a wavelength-dependent multiplicative function such that the SED exactly reproduces the observed DES $griz$ photometry (often referred to as ‘mangling’; Hsiao et al., 2007; Conley et al., 2008).
@@ 2.1 SN Ia distances and bias corrections
where $m_{x}$, $x_{1}$ and $c$ are the SALT3 SN light-curve parameters, $\alpha$ and $\beta$ are global parameters parametrizing the stretch–luminosity and colour–luminosity relations, and M is the absolute magnitude of a SN Ia with $x_{1}=0$ and $c=0$. $\mu_{\mathrm{bias}}$ is the bias term that corrects for various observational and astrophysical selection effects.
@@
where $\mu_{\mathrm{obs}}$ is defined in equation 1 and $\mu_{\mathrm{model}}$ is the distance modulus in a reference cosmology. The Hubble residuals indicate whether the measured distance to a given SN Ia is larger or smaller than that expected in the reference cosmology at the redshift of that SN; negative residuals indicate smaller distances.
In our analysis we differ from the main DES-SN5YR analysis in three ways. Firstly, we do not include a ‘$\gamma G_{\mathrm{host}}$’ term in equation 1. When used, such a term is designed to encapsulate any residual dependencies between the distances estimated to SNe Ia and their host galaxy properties (such as a stellar mass step) that are not corrected for in the light-curve fitting or bias correction steps. As a goal of our paper is to investigate such host galaxy effects, we do not apply this correction.
@@
In estimating our distances using BBC 1D, we recover parameter values of $\alpha=0.159\pm 0.009$ and $\beta=2.73\pm 0.05$. These differ from those reported in table 5 of Vincenzi et al. (2024), as those values were estimated using the BBC 4D bias corrections, but are similar to other estimates of $\alpha$ and $\beta$ found for the DES-SN5YR sample using 1D corrections (e.g., Kelsey et al., 2023).
@@ 2.2 SN–host separations
SN celestial coordinates ($x_{\mathrm{SN}}$, $y_{\mathrm{SN}}$) were measured by DES using point-spread function (PSF) fitting photometry (Kessler et al., 2015), with the coordinates computed using the weighted-average from each measurement during an observing ‘season’. Host galaxies were detected and measured in the DES image stacks by (Wiseman et al., 2020) using Source Extractor. We use the celestial coordinates (image centroids) of the Source Extractor detections ($x_{\mathrm{host}}$, $y_{\mathrm{host}}$) measured from images with the same world coordinate system as the images on which $x_{\mathrm{SN}}$, $y_{\mathrm{SN}}$ are determined. The statistical uncertainties in both ($x_{\mathrm{SN}}$, $y_{\mathrm{SN}}$) and ($x_{\mathrm{host}}$, $y_{\mathrm{host}}$) are small: the statistical uncertainties from the PSF fitting of ($x_{\mathrm{SN}}$, $y_{\mathrm{SN}}$) and the statistical uncertainties reported by Source Extractor for ($x_{\mathrm{host}}$, $y_{\mathrm{host}}$) have a median uncertainty of 8 mas or 0.03 pixels. The overall DECam astrometric solution is accurate to 3–6 milliarcseconds or $\simeq$0.02 pixels (Bernstein et al., 2017); it is critical that this number is small as the SN and host positions are not measured from the same images and thus represents the systematic uncertainty. All these positional uncertainties are very small and we neglect them in the remainder of our analysis.
These coordinates then define the angular separation (or projected galactocentric distance), $\Delta\theta$, between the SN and the centre of the host galaxy, i.e.,
@@
To measure DLRs, we use the Source Extractor basic shape parameters A (semi-major axis) and B (semi-minor axis) that define the elliptical shape of the deblended galaxy. The scales of A and B are set by the second-order moments (variance) of the galaxy’s profile along the A and B axes (e.g., Stobie, 1980). In particular, we do not use isophotal apertures, where if the same galaxy were observed at increasing redshift, the DLR would become smaller due to cosmological surface brightness dimming (see discussion in, e.g., Infante, 1987). The Kron-like apertures used here contain a consistent 94 per cent of the galaxy light (Bertin & Arnouts, 1996).
@@ 3.2.1 Hubble residuals as a function of $d_{\mathrm{DLR}}$
The trend itself is not unexpected given the known relationships between Hubble residual and stellar mass (e.g., left panel of Fig. 4): regions within $d_{\mathrm{DLR}}<0.5$ are likely to be more similar to the older, passive stellar environments of massive galaxies where more negative Hubble residuals are also observed. A logical next step is to test how the Hubble residual–$d_{\mathrm{DLR}}$ trend relates to the well-known stellar mass and host galaxy colour steps.
@@ 3.2.2 Host stellar mass step and $d_{\mathrm{DLR}}$
We show the host galaxy stellar mass step separately for SNe Ia in bins of $d_{\mathrm{DLR}}$ and find that the size of the stellar mass step does depend on the galactocentric distance (Fig. 5). We split the sample at $d_{\mathrm{DLR}}=1$: within that radius, SNe have a mass step of $0.100\pm 0.014$ mag ($6.9\,\sigma$), whereas for $d_{\mathrm{DLR}}>1$ the step is consistent with zero ($0.036\pm 0.018$ mag; $2.0\,\sigma$). The difference in step size between $d_{\mathrm{DLR}}>1$ and $d_{\mathrm{DLR}}\leq 1$ is $0.064\pm 0.023$ mag. Step sizes and their significance are reported in Table 2. At this point, we make no attempt to optimise the step size difference between inner and outer regions, and selected $d_{\mathrm{DLR}}=1$ only for its convenient integer nature.
@@
Sample Number Mean $\Delta\mu$ in Mean $\Delta\mu$ in Size of the Significance r.m.s. in r.m.s. in
(LM hosts, HM hosts) low-mass hosts high-mass hosts mass step of step low-mass hosts high-mass hosts
@@
$d_{\mathrm{DLR}}\leq 1$ 1028 (310, 718) $0.065\pm 0.012$ $-0.035\pm 0.009$ $0.100\pm 0.014$ $6.9\sigma$ 0.205 0.232
$d_{\mathrm{DLR}}>1$ 505 (162, 343) $0.023\pm 0.014$ $-0.014\pm 0.011$ $0.036\pm 0.018$ $2.0\sigma$ 0.180 0.212
$z<0.6$ 982 (329, 653) $0.049\pm 0.009$ $-0.034\pm 0.008$ $0.083\pm 0.012$ $6.9\sigma$ 0.166 0.201
$d_{\mathrm{DLR}}\leq 1$ 652 (220, 432) $0.063\pm 0.012$ $-0.041\pm 0.010$ $0.104\pm 0.015$ $6.7\sigma$ 0.174 0.211
$d_{\mathrm{DLR}}>1$ 330 (109, 221) $0.021\pm 0.014$ $-0.022\pm 0.012$ $0.043\pm 0.018$ $2.4\sigma$ 0.141 0.178
$P_{\mathrm{Ia}}>0.9$ 1442 (449,991) $0.050\pm 0.009$ $-0.028\pm 0.007$ $0.079\pm 0.012$ $6.8\sigma$ 0.195 0.222
$d_{\mathrm{DLR}}\leq 1$ 968 (295, 673) $0.065\pm 0.012$ $-0.034\pm 0.009$ $0.100\pm 0.015$ $6.8\sigma$ 0.203 0.227
$d_{\mathrm{DLR}}>1$ 472 (154, 318) $0.021\pm 0.014$ $-0.018\pm 0.012$ $0.039\pm 0.018$ $2.1\sigma$ 0.174 0.212
@@
We perform two additional tests of the change in mass step: we restrict our sample to events at $z<0.6$ where selection effects are smaller, and we select events with $P_{\mathrm{Ia}}>0.9$ (instead of 0.5; see Section 2) to remove those events least likely to be SNe Ia. The results are in Table 2. The difference in step sizes between $d_{\mathrm{DLR}}>1$ and $d_{\mathrm{DLR}}\leq 1$ for these two samples is $0.061\pm 0.023$ mag in both cases, compared to $0.064\pm 0.023$ mag in our full sample, an almost identical result.
To determine if the observed difference in mass step size between inner and outer regions have sensitivity to the choice of $d_{\mathrm{DLR}}=1$, we use smaller $d_{\mathrm{DLR}}$ bins, each with a similar number of SNe Ia. We also fit a straight line to the data. The results are shown in Fig. 7. The figure shows a clear difference in step size between $d_{\mathrm{DLR}}\leq 1$ and $d_{\mathrm{DLR}}>1$, i.e., the result in Fig. 5 is not dependent on the choice of $d_{\mathrm{DLR}}=1$ to divide the sample.
@@
Figure 8: As Fig. 5 but for host galaxy rest-frame $U-R$ colour. The step sizes are $0.107\pm 0.013$ mag (left) and $0.029\pm 0.018$ mag (right).